If 3/4 ib. of candy cost $20.50, how much would 1lb. of candy cost.
Answer by TimothyLamb(4379)

c/w = 20.50/(3/4)
c/1 = 20.50/(3/4)
c = 1*20.50/(3/4)
c = 27.3333333333
---
answer:
1 lb costs $27.33
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---